OpenStudy (anonymous):

Mr. Evans paid $455 for a 7-day car rental. At that rate, how much would he pay for a 10-day rental? Options- A. $65 B. $650 C. $465 D. $462

OpenStudy (anonymous):

@iGreen @babygirl121314 @Harsha19111999

OpenStudy (anonymous):

455 / 7 = 65

OpenStudy (anonymous):

devide 7/455 then multiply ur answer x 10

OpenStudy (anonymous):

so he earned $65 a day

OpenStudy (harsha19111999):

For 7 days it is $455. Then for 1 day what is it? 455/7 = 65. And for 10 days it is 10*65 =$650

OpenStudy (anonymous):

can you complete that?

OpenStudy (anonymous):

650$$ for 10 days

OpenStudy (anonymous):

B

OpenStudy (anonymous):

650 is my answer... is that correct?

OpenStudy (anonymous):

yes 100% sure! ur in 6th grade math?

OpenStudy (anonymous):

No :)

OpenStudy (anonymous):

5th?

OpenStudy (anonymous):

No :D

OpenStudy (anonymous):

4th?

OpenStudy (harsha19111999):

Yes 650 is correct

OpenStudy (anonymous):

NO!! :D @qdbug

OpenStudy (anonymous):

how old r u?

OpenStudy (anonymous):

Do not ask personal questions

OpenStudy (anonymous):

i promise im not a moder

OpenStudy (anonymous):

im katie and im 13

OpenStudy (anonymous):

Oh Hi!